I need help setting this up.
Peanuts cost $3 per pound, almonds $4 per pound, and cashews $8 per pound. An event planner wishes to purchase 140 pounds of mixed nuts for $6. per pound, in which there are twice as many peanuts as almond. Can you advise the event planner as to how many of each kind of nut should be in the mixture?

Answer by mananth(12270)


Let the quantity of peanuts to be mixed =x
Almonds will be 2x
Let the cashew nuts mixed be y
3x+4*2x+8y=140*6=840
11x+8y=840 ---------1
3x+y= 140------------2
Multiply eq2 by 8 and subtract from 1
24x+8y=1120

11x+8y=840
24x+8y=1120
-13x= -280
X=280/13
=21.54 = peanuts
2x= 43.08 =almonds
Y=75.38 = cashew nuts
M Ananth------------- mananth@hotmail.com